# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 4.1. Datasets
The most of datasets used in the credibility assessment works are annotated at the content level (see Section 2.2 for the overview of levels of credibility assessment). Table 2 provides an overview of the selected datasets created or adapted for the purpose of credibility assessment task.
@@
El Ballouli et al. (2017) collected 17 million tweets in a period of two weeks that were written in Arabic and contained any hashtag. After preprocessing and grouping tweets by their hashtags, a topic-independent subset of 9,000 tweets was selected. Each of these tweets was annotated by three annotators on a binary scale (credible, non-credible), while the majority vote was used to determine the final label.
Due to a lack of sufficiently large credibility-annotated datasets, some authors (e.g., (Wu et al., 2020; Podgurski et al., 2022; Leite et al., 2024)) decided to use available fake news datasets, such as LIAR (Wang, 2017), Weibo10, FakeNewsNet (Shu et al., 2020), FakeNewsAMT (Pérez-Rosas et al., 2018) or Celebrity dataset (Pérez-Rosas et al., 2018). In such a case, various fake-news labels are mapped to credibility labels by taking an assumption that a fake news content is considered to be non-credible. Qureshi and Malick (2023) also proceeded from the FakeNewsNet dataset, however, only to identify a potential set of tweets that were further manually annotated by 12 experts for credibility on a 5-point scale. The resulting dataset consists of 4,958 tweets, out of them some were discarded due to low annotators’ agreement.
Another group of datasets used in the existing works relates to credibility annotated at source (webpage) level. Microsoft Credibility (Schwarz and Morris, 2011) dataset consist of top 40 search results on 25 pre-defined queries on the topics of Health, Politics, Finance, Environmental Science, and Celebrity News, which were manually annotated for credibility on a 5-point scale. Another dataset, Content Credibility Corpus (C3) (Kakol et al., 2017), consists of 15,750 evaluations of 5,543 pages by more than 2,000 annotators recruited through Amazon Mechanical Turk.
@@
Dataset | Language | # Instances | Content type | Classes | Used by | |||||
@@
Weibo101010https://service.account.weibo.com | Chinese | 18,000 | Microblogs |
@@ 4.2. Methods and models
The methods and models used in the research works falling into this group typically employ a methodology consisting of three steps. Firstly, a set of credibility signals is selected. Secondly, various techniques are employed for their automatic detection. Finally, detected credibility signals are then either used to predict an overall credibility which can be subsequently utilized for example as an input to information retrieval engine. In the following description, we provide an in-depth description of each of these three steps.
Credibility signals selection. In the earliest works, credibility signals were rather approximated by various (mostly shallow) linguistic features that can be easily detected by the machine, such as a number of words or exclamation marks. While being potentially helpful content representations for traditional machine learning algorithms, people would not usually use such features for credibility assessment due to the absence of their direct interpretability. Weerkamp and De Rijke (2008) adopted a subset of 11 such features for blog posts proposed in (Rubin and Liddy, 2006) while focusing on those that are text-based and can be easily detected automatically, such as capitalization, shouting, spelling or post length. Similarly, Kang et al. (2012) opted for 12 numeric and 7 binary content-based credibility indicators for Twitter, such as a positive sentiment factor or a number of mentions. Approximately half of these indicators were adopted from the previous work by Castillo et al. (2011).
@@
In contrast to works mentioned so far, the following works selected signals from the list created by the Credible Web Community Group (see Section 2.2). At first, Podgurski et al. (2022) proposed a modular credibility assessment system of web pages using the 23 most frequent content- and context-based signals. The employed signals range from naive numerical and linguistic features, such as the number of external links, linguistic statistics (average text and word count, overall word count, etc.), spelling and grammar, domain endings and author information, to more complex features, such as sentiment, readability and presence of a clickbait title. Similarly, Leite et al. (2024) employed 18 credibility signals, including advanced ones that have not been considered in the previous works, such as call to action, impoliteness, sensationalism or explicitly unverified claims.
@@
By proceeding from such superficial linguistic features to more advanced credibility signals, also their detection methods became more complex. For sentiment analysis various lexicon and rule-based sentiment analysis were employed, such as Vader for English (Esteves et al., 2018; Podgurski et al., 2022), or ArSenL for Arabic (El Ballouli et al., 2017). For text category classification, well-known approaches employed in other NLP tasks were adopted, such binary multinomial Naive Bayes (NB) classifiers or Latent Semantic Analysis (LSA) (Esteves et al., 2018). Various NLP libraries were used for automatic detection of additional credibility signals, such as LanguageTool library for detection of grammar & spelling errors or TextBlob for subjectivity detection (Podgurski et al., 2022). Even further, Qureshi and Malick (2023) detected the selected credibility signals by means of the classification methods proposed in the previous research works.
Surprisingly, adoption of LLMs for detection of multiple categories of credibility signals is only very rare despite a great potential of LLMs to target a wide range of various credibility signals categories. To this end, Leite et al. (2024) used 3 instruction-tuned LLMs (GPT-3.5-Turbo, Alpaca-LoRA-30B and OpenAssistant-LLaMa-30B) with a specific prompt designed for each credibility signals (e.g., ”Does the article make use of sensationalist claims?”).
Credibility score prediction. To estimate overall credibility from the credibility signals, Weerkamp and De Rijke (2008) opted for simple weighting schemata. The resulting credibility score was subsequently incorporated into a blog post retrieval model. Kang et al. (2012) proposed three computational models (based on a weighted combination, and on a probabilistic language-based approach) for assessing tweet’s credibility, using not only content-based but also social and hybrid strategies. Experiments on the test set consisting of 1,023 instances (while using 10-fold cross-validation) revealed that social model was able to outperform content-based and hybrid model. This can be explained by a short textual content of tweets, as well as very shallow linguistic and possibly noisy credibility signals, such as a presence of exclamation mark. Podgurski et al. (2022) combined individual signals’ subscores with signal weights through a linear combination function. Signal weights were determined by authors empirically by considering prior research results, signal measurement accuracy, and experimental calculations on test data.
@@
For the purpose of feature selection, Qureshi and Malick (2023) utilized the Kolmogorov–Smirnov (KS) test to identify discriminating features. Subsequently, a wide range of machine learning algorithms (12 regression and 10 classification ones) were used to predict original 5-scale as well as simplified binary credibility score. Results revealed that context-bases signals, such as high user influence and medium-to-high spread count, are very indicative of credible tweets. Credible tweets are normally spread by topic experts (the topic of the tweet is in the top 3 topics for that user), typically do not contain deceptive information and are not posted by malicious accounts. In terms of content-based signals, credible tweets have high informativeness and tent to not contain hate speech or links to low-credible sources.
@@
Finally, Leite et al. (2024) used Prompted Weak Supervision (PWS) on the top of LLM-predicted credibility signals. This approach was compared with unsupervised and supervised baselines. In the supervised finetuning scenario, a RoBERTa-Base model was finetuned with the ground-truth labels. In the zero-shot scenario, an instruction-tuned LLM based on LLaMa2 was prompted without any signals or in-context examples. Experimental results on four datasets (FakeNewsNet - GossipCop, FakeNewsNet - PolitiFact, FakeNewsAMT, and Celebrity) showed that Prompted Weak Supervision outperformed the zero-shot baseline by , and achieved of the performance of the state-of-the-art supervised baseline. Furthermore, in cross-domain settings, where the domain of the train set differs from the domain of the test set (e.g., Politics and Gossip), Prompted Weak Supervision outperformed the supervised baseline by . Lastly, the authors study the association between credibility signals and veracity through (i) a statistical test, where out of the signals were shown to have an association with veracity, and (ii) an ablation study, in which signals were individually removed from the train set in order to inspect their contribution to the model’s performance. This showed that the method’s strength is in the combination of a wide range of signals rather than relying on a small set of signals that could be strong predictors of veracity. Through these analyses, the authors verified that some credibility signals are domain-specific. For example, signals such as Source Credibility and Misleading about Content improve performance mostly for the political domain, while others such as Expert Citation and Call to Action show benefits in entertainment news. In addition to the efficiency of this approach in terms of being independent of the costly annotation of long news articles with each signal, the main advantage of this method is generalisation capability. The authors found that their approach outperforms the fine-tuned methods on the out-of-domain test data. This result is particularly important in the modern era of misinformation, when new fake news emerge every day, making it impossible to have up-to-date annotated data.
@@ 4.3. Tools and services
While the systematic paper review have not revealed any specific demo papers describing tools or services dedicated to credibility assessment, various publicly available solutions exist. Many of them are on a border with false information detection. From 2019, the Credibility Coalition maintain a CredCatalog111111https://credibilitycoalition.org/credcatalog/ – a catalog of initiatives that have a stated aim to improve information quality. Besides various organizations (like fact-checking or academic instititions), it provides overview of several relevant tools.
Besides that, there are commercial projects aiming to help end-users to evaluate credibility of the online content directly in their browsers. NewsGuard121212https://www.newsguardtech.com/ provides News Reliability Ratings for news outlets based on nonpartisan journalistic criteria. Similarly, GroundNews131313https://ground.news/ rate news stories according to their bias distribution (left-/right-leaning bias) or factuality. The labelling is, however, done rather indirectly through sources writing about such news stories, instead of analysing a content of individual news articles themselves.
Finally, an automatic analysis of several credibility signals (persuasion techniques, subjectivity, or machine-generated text among others) is available as a part of the Assistant tool in the well-known Verification plugin141414https://www.veraai.eu/category/verification-plugin developed and further enhanced as a part of EU-funded projects (InVID, WeVerify and vera.ai). The Assistant tool allows users to provide an URL or a local file as input, from which the text is extracted and various text analysis are performed, credibility signals being a recent addition to them.
@@ 4.4. Discussion
Limited research focus with absence of LLM-based solutions. As already shown in Figure 3, the number of works explicitly addressing credibility assessment by taking multiple categories of credibility signals into account is considerably lower in comparison with other individual signal categories. This is in contrast to fake-news detection, which attracted a plethora of research (and also practitioners’ and public) attention.
In parallel, our literature survey pointed out a surprising lack of approaches utilizing LLMs. At the same time, LLMs provide a great opportunity to automatize detection of multiple credibility signals even in a zero-shot settings, as the work by Leite et al. (2024) clearly demonstrated. Even further improvement in the terms of accuracy can be achieved by employing in-context learning, or instruction-tuning LLMs. While such LLM-based approaches would inquire higher computational costs, existing Parameter-efficient Fine Tuning (PEFT) techniques (Xu et al., 2023) may employed if necessary.
Lack of multilingual datasets with annotated credibility signals. Similarly to a limited research focus, also the situation with credibility-annotated datasets is falling behind fake-news research. There are very few datasets with manual (human) annotations at the level of individual credibility signals as well as overall credibility (Zhang et al., 2018; Bhuiyan et al., 2020). These datasets are, unfortunately, very small preventing employment of language models even in few-shot settings. Despite the declared original intentions to extend them in future, to the best of our knowledge, no follow-up datasets have been created so far.
The remaining datasets are larger (containing hundreds or thousands of instances), nevertheless, annotated only at the level of overall credibility (see Table 2). Absence of datasets providing annotations of multiple categories of credibility signals for the same set of instances, prevents multi-task training of models. Multi-task approaches may provide an interesting potential, especially since many credibility signals inherently share some underlying similarities that can contribute to a better performance (as the results of Prompted Weak Supervision (PWS) already demonstrated in (Leite et al., 2024)).
@@ 5. Factuality, subjectivity and bias
Factuality is a signal reflecting a degree of certainty regarding the possibility of events. The literature on factuality traditionally distinguishes between individual event factuality detection (EFD) (Li et al., 2020b) and document-level event factuality identification (DEFI) (Cao et al., 2021) tasks. The main distinction between these two sub-tasks is in the scope of information concerned – while the EFD task considers individual event mentions, DEFI amis to aggregate several mentions of a certain event within a document.
@@
Finally, bias represents a similar but more complex phenomenon of imbalance in terms of opinions or facts. As highlighted by (Nakov et al., 2021b), there is no one single concept of bias among scholars. However, in many cases bias is seen as a systematic favouring of a certain ideology when covering information (Waldman and Devitt, 1998). This can be expressed in deliberately withholding certain part of information that contradict the favourable point of view (Smith et al., 2001) or vice versa, specifically searching for the facts that cover information from a certain political or ideological viewpoint (Hassell et al., 2020). In some cases, even if the choice of information sources is unbiased, the presentation of information in those sources may be performed in a biased manner that highlights the importance of only certain facts. Some instances of biased presentation include framing (Piskorski et al., 2023; Entman, 2007), opinionated reporting style (Piskorski et al., 2023; Soroka, 2012) and even visual clues (Barrett and Barrington, 2005) used to distort the perception of information.

### v2

@@ 4.1. Datasets
Most of datasets used in the credibility assessment works are annotated at the content level (see Section 2.2 for the overview of levels of credibility assessment). Table 2 provides an overview of the selected datasets created or adapted for the purpose of credibility assessment task.
@@
El Ballouli et al. (2017) collected 17 million tweets in a period of two weeks that were written in Arabic and contained at least one hashtag. After preprocessing and grouping tweets by their hashtags, a topic-independent subset of 9,000 tweets was selected. Each of these tweets was annotated by three annotators on a binary scale (credible, non-credible), while the majority vote was used to determine the final label.
Due to a lack of sufficiently large credibility-annotated datasets, some authors (e.g., (Wu et al., 2020; Podgurski et al., 2022; Leite et al., 2024)) decided to use available false information (fake news) datasets, such as LIAR (Wang, 2017), Weibo (Wu et al., 2020), FakeNewsNet (Shu et al., 2020), FakeNewsAMT (Pérez-Rosas et al., 2018) or Celebrity dataset (Pérez-Rosas et al., 2018). In such a case, various false information labels are mapped to credibility labels by taking an assumption that a false content is considered to be non-credible. Qureshi and Malick (2023) also proceeded from the FakeNewsNet dataset, however, only to identify a potential set of tweets that were further manually annotated by 12 experts for credibility on a 5-point scale. The resulting dataset consists of 4,958 tweets, out of them some were discarded due to low annotators’ agreement.
Another group of datasets used in the existing works relates to credibility annotated at a source (webpage) level. Microsoft Credibility (Schwarz and Morris, 2011) dataset consists of top 40 search results on 25 pre-defined queries on the topics of Health, Politics, Finance, Environmental Science, and Celebrity News, which were manually annotated for credibility on a 5-point scale. Another dataset, Content Credibility Corpus (C3) (Kakol et al., 2017), consists of 15,750 evaluations of 5,543 pages by more than 2,000 annotators recruited through Amazon Mechanical Turk. The textual credibility signals annotated in this dataset include readability, language quality, informativeness, completeness, and objectivity.
@@
| Dataset | Language(s) | # Instances | Content type | Classes | Used by | |||||
@@
| Weibo (Wu et al., 2020) | Chinese | 18,000 | Microblogs |
@@ 4.2. Methods and models
The methods and models used in the research works falling into this group typically employ a methodology consisting of three steps. Firstly, a set of credibility signals is selected. Secondly, various techniques are employed for their automatic detection. Finally, detected credibility signals are then used to predict an overall credibility which can be subsequently utilized for example as an input to information retrieval engine. In the following description, we provide an in-depth description of each of these three steps.
Credibility signals selection. In the earliest works, credibility signals were rather approximated by various (mostly shallow) linguistic features that can be easily detected by the machine, such as a number of words or exclamation marks. While being potentially helpful for traditional machine learning algorithms, people would not usually use such features for credibility assessment due to the absence of their direct interpretability. Weerkamp and De Rijke (2008) adopted a subset of 11 such features for blog posts proposed in (Rubin and Liddy, 2006) while focusing on those that are text-based and can be easily detected automatically, such as capitalization, shouting, spelling or post length. Similarly, Kang et al. (2012) opted for 12 numeric and 7 binary content-based credibility indicators for Twitter, such as a positive sentiment factor or a number of mentions. Approximately half of these indicators were adopted from the previous work by Castillo et al. (2011).
@@
In contrast to works mentioned so far, the following works selected signals from the list created by the Credible Web Community Group (see Section 2.2). At first, Podgurski et al. (2022) proposed a modular credibility assessment system of web pages using the 23 most frequent content- and context-based signals. The employed signals range from naive numerical and linguistic features, such as a number of external links, linguistic statistics (average text and word count, overall word count, etc.), spelling and grammar, domain endings and author information, to more complex features, such as sentiment, readability and presence of a clickbait title. Similarly, Leite et al. (2024) employed 18 credibility signals, including advanced ones that have not been considered in the previous works, such as call to action, impoliteness, sensationalism or explicitly unverified claims.
@@
By proceeding from such superficial linguistic features to more advanced credibility signals, also their detection methods became more complex. For sentiment analysis various lexicon and rule-based sentiment analysis were employed, such as Vader for English (Esteves et al., 2018; Podgurski et al., 2022), or ArSenL for Arabic (El Ballouli et al., 2017). For text category classification, well-known approaches employed in other NLP tasks were adopted, such binary multinomial Naïve Bayes (NB) classifiers or Latent Semantic Analysis (LSA) (Esteves et al., 2018). Various NLP libraries were used for automatic detection of additional credibility signals, such as LanguageTool library for detection of grammar & spelling errors or TextBlob for subjectivity detection (Podgurski et al., 2022). Even further, Qureshi and Malick (2023) detected the selected credibility signals by means of the classification methods proposed in the previous research works.
Surprisingly, adoption of LLMs for detection of credibility signals is only very rare despite a great potential of LLMs to target a wide range of various credibility signals categories. To this end, Leite et al. (2024) used 3 instruction-tuned LLMs (GPT-3.5-Turbo, Alpaca-LoRA-30B and OpenAssistant-LLaMa-30B) with a specific prompt designed for each credibility signals (e.g., “Does the article make use of sensationalist claims?”).
Credibility label/score prediction. To estimate overall credibility from the credibility signals, Weerkamp and De Rijke (2008) opted for simple weighting schemata. The resulting credibility score was subsequently incorporated into a blog post retrieval model. Kang et al. (2012) proposed three computational models (based on a weighted combination, and on a probabilistic language-based approach) for assessing tweet’s credibility, using not only content-based but also social and hybrid strategies. Experiments on the test set consisting of 1,023 instances (while using 10-fold cross-validation) revealed that social model was able to outperform content-based and hybrid model. This can be explained by a short textual content of tweets, as well as very shallow linguistic and possibly noisy credibility signals, such as a presence of exclamation mark. Podgurski et al. (2022) combined individual signals’ subscores with signal weights through a linear combination function. Signal weights were determined by authors empirically by considering prior research results, signal measurement accuracy, and experimental calculations on test data.
@@
For the purpose of feature selection, Qureshi and Malick (2023) utilized the Kolmogorov–Smirnov (KS) test to identify discriminating features. Subsequently, a wide range of machine learning algorithms (12 regression and 10 classification ones) were used to predict original 5-scale as well as simplified binary credibility label. Results revealed that context-bases signals, such as high user influence and medium-to-high spread count, are very indicative of credible tweets. Credible tweets are normally spread by topic experts (the topic of the tweet is in the top 3 topics for that user), typically do not contain deceptive information and are not posted by malicious accounts. In terms of content-based signals, credible tweets have high informativeness and tent to not contain hate speech or links to low-credible sources.
@@
Finally, Leite et al. (2024) used Prompted Weak Supervision (PWS) on the top of LLM-predicted credibility signals. This approach was compared with unsupervised and supervised baselines. In the supervised finetuning scenario, a RoBERTa-Base model was finetuned with the ground-truth labels. In the zero-shot scenario, an instruction-tuned LLM based on LLaMa2 was prompted without any signals or in-context examples. Experimental results on four datasets (FakeNewsNet - GossipCop, FakeNewsNet - PolitiFact, FakeNewsAMT, and Celebrity) showed that Prompted Weak Supervision outperformed the zero-shot baseline by , and achieved of the performance of the state-of-the-art supervised baseline. Furthermore, in cross-domain settings, where the domain of the train set differs from the domain of the test set (e.g., Politics and Gossip), Prompted Weak Supervision outperformed the supervised baseline by . Lastly, the authors study the association between credibility signals and veracity through (i) a statistical test, where out of the signals were shown to have an association with veracity, and (ii) an ablation study, in which signals were individually removed from the train set in order to inspect their contribution to the model’s performance. This showed that the method’s strength is in the combination of a wide range of signals rather than relying on a small set of signals that could be strong predictors of veracity. Through these analyses, the authors verified that some credibility signals are domain-specific. For example, signals such as Source Credibility and Misleading about Content improve performance mostly for the political domain, while others such as Expert Citation and Call to Action show benefits in entertainment news. In addition to the efficiency of this approach in terms of being independent of the costly annotation of long news articles with each signal, the main advantage of this method is generalisation capability. The authors found that their approach outperforms the fine-tuned methods on the out-of-domain test data. This result is particularly important in the constantly-evolving online environment, when new non-credible and false information emerge every day, making it impossible to have up-to-date annotated data.
@@ 4.3. Tools and services
While several services dedicated to credibility assessment exist, many of them are on a border with false information detection. From 2019, the Credibility Coalition maintain a CredCatalog999https://credibilitycoalition.org/credcatalog/ – a catalogue of initiatives that have a stated aim to improve information quality. Besides various organizations (like fact-checking or academic institutions), it provides overview of relevant tools.
Besides that, there are commercial projects supporting end-users in evaluating the credibility of the online content directly in their browsers. NewsGuard101010https://www.newsguardtech.com/ provides News Reliability Ratings for news outlets based on nonpartisan journalistic criteria. Similarly, GroundNews111111https://ground.news/ rate news stories according to their bias distribution (left-/right-leaning bias) or factuality (in the meaning of source reliability and general factual accuracy). The labelling is, however, done rather indirectly through sources writing about such news stories, instead of analysing a content of individual news articles themselves.
Another relevant tool is the Tanbih (Zhang et al., 2019) – a news aggregator that performs media-level and article-level analyses, aiming to help users better understand the content they consume. It organizes news articles into event-based clusters and generates media outlet profiles, which include indicators such as the general factuality (in the meaning of factual accuracy) of reporting, the degree of propagandistic content, hyper-partisanship, leading political ideology, overall framing, and the stance of the outlet toward various claims and topics.
Finally, an automatic analysis of several credibility signals (persuasion techniques, subjectivity, or machine-generated text among others) is available as a part of the Assistant tool in the well-known Verification plugin121212https://www.veraai.eu/category/verification-plugin developed and further enhanced as a part of EU-funded projects (InVID, WeVerify and vera.ai). The Assistant tool allows users to provide a URL or a local file as input, from which the text is extracted and various text analysis are performed, credibility signals being a recent addition to them.
@@ 4.4. Discussion
Limited research focus with absence of LLM-based solutions. As already shown in Figure 5, the number of works explicitly addressing automatic credibility assessment is considerably lower in comparison with other individual signal categories. This is in contrast to false information detection, which attracted a plethora of research (and also practitioners’ and public) attention.
In parallel, our literature survey pointed out a lack of approaches utilizing LLMs. At the same time, LLMs provide a great opportunity to automatize detection of multiple credibility signals even in a zero-shot settings, as the work by Leite et al. (2024) clearly demonstrated. Even further improvement in the terms of accuracy can be achieved by employing in-context learning, or instruction-tuning LLMs. While such LLM-based approaches would inquire higher computational costs, existing Parameter-efficient Fine Tuning (PEFT) techniques (Xu et al., 2023) may be employed if necessary.
Lack of multilingual datasets with annotated credibility signals. Similarly to a limited research focus, also the situation with credibility-annotated datasets is falling behind false information research. There are very few datasets with manual (human) annotations at the level of individual credibility signals as well as overall credibility (Zhang et al., 2018; Bhuiyan et al., 2020). Unfortunately, these datasets are very small, which prevents the use of language models even in few-shot settings. Despite the declared original intentions to extend them in future, to the best of our knowledge, no follow-up datasets have been created so far.
The remaining datasets are larger (containing hundreds or thousands of instances), nevertheless, annotated only at the level of overall credibility (see Table 2). Absence of datasets providing annotations of multiple categories of credibility signals for the same set of instances, prevents multi-task training of models. Multi-task approaches hold considerable potential, particularly because many credibility signals inherently share underlying similarities that can enhance model performance, as already demonstrated by the results of Prompted Weak Supervision (PWS) in (Leite et al., 2024).
@@ 5. Factuality, subjectivity and bias
Within false information research, the term factuality has been used to refer to a variety of related problems, with no unified ontology to define its levels. For example, Nakov et al. (2024) equate factuality with factual accuracy (i.e., veracity) and propose a four-level ontology encompassing the claim, article, user, and source factuality levels. In this interpretation, factuality is seen not as as a credibility signal, but as the final verdict regarding truthfulness of information. In the context of large language models (LLMs), factuality is typically defined as a model’s ability to generate outputs grounded in established real-world facts (Wang et al., 2023b). This aligns with the Cambridge English Dictionary’s definition of factual as “based on or containing facts”131313https://dictionary.cambridge.org/dictionary/english/factual. From an NLP perspective, however, factuality is often approached differently. Rather than relying solely on external fact-checking, NLP literature tends to treat factuality as a contextual signal – reflecting the degree of certainty a speaker conveys about the occurrence of events (Li et al., 2020b; Cao et al., 2021; Qian et al., 2019; Li et al., 2024c; Rovera et al., 2025). As identified through the systematic literature review conducted as part of this work, factuality in this view functions as a credibility signal rather than an overall veracity assessment, expressing how confidently an event is presented as having occurred or not occurred. This more fine-grained perspective follows earlier work by Saurí and Pustejovsky (2009), which defines factuality as the extent to which events are portrayed as corresponding to real-world situations, hypothetical scenarios, or uncertain interpretations. Under this definition, the task of factuality detection can be used as the first step towards identifying check-worthy claims (discussed in Section 7), as only the events that the speaker is confident about typically form claims. In the rest of this section, we use the terms factuality and event factuality interchangeably. Where the meaning is not obvious from the context or could lead to ambiguity, we distinguish whether event factuality, factual accuracy or a different notion of factuality is meant there.
The literature on event factuality traditionally distinguishes between individual event factuality detection (EFD) (Li et al., 2020b) and document-level event factuality identification (DEFI) (Cao et al., 2021) tasks. The main distinction between these two sub-tasks is in the scope of information concerned – while the EFD task considers individual event mentions, DEFI aims to aggregate several mentions of a certain event within a document.
@@
Finally, bias represents a similar but more complex phenomenon of imbalance in terms of opinions or facts. As highlighted by (Nakov et al., 2024), there is no one single concept of bias among scholars. However, in many cases bias is seen as a systematic favouring of a certain ideology when covering information (Waldman and Devitt, 1998). This can be expressed in deliberately withholding certain part of information that contradict the favourable point of view (Smith et al., 2001) or vice versa, specifically searching for the facts that cover information from a certain political or ideological viewpoint (Hassell et al., 2020). In some cases, even if the choice of information sources is unbiased, the presentation of information in those sources may be performed in a biased manner that highlights the importance of only certain facts. Some instances of biased presentation include framing (Piskorski et al., 2023; Entman, 2007), opinionated reporting style (Piskorski et al., 2023; Soroka, 2012) and even visual clues (Barrett and Barrington, 2005) used to distort the perception of information.